# ProbTS: Benchmarking Point and Distributional Forecasting across Diverse Prediction Horizons

### v4

@@ D.1 Impact of Data Scale
D.2 Statistical and Gradient Boosting Decision Tree Baselines
@@
D.3 Experiments on Univariate Datasets
@@
Table 16 presents experimental results for representative methods, consistent with our initial observations. Probabilistic estimation methods like GRU NVP and TimeGrad excel on datasets with complex distributions (e.g., M4-Weekly and M5), while simpler point forecasting methods such as DLinear and PatchTST perform well on datasets with relatively simple data distribution, like TOURISM-Monthly. Both autoregressive and non-autoregressive decoding schemes show comparable performance in short-term forecasting, as discussed in the main paper."
@@
D.4 Experiments on Synthetic Datasets
@@ D.6 Model Efficiency
D.7 Further Discussion on Cross-channel Interactions

### v5

@@ D.2 The Impact of Data Scale
D.3 Statistical and Gradient Boosting Decision Tree Baselines
@@
D.4 Experiments on Univariate Datasets
@@
Table 21 presents experimental results for representative methods, consistent with our initial observations. Probabilistic estimation methods like GRU NVP and TimeGrad excel on datasets with complex distributions (e.g., M4-Weekly and M5), while simpler point forecasting methods such as DLinear and PatchTST perform well on datasets with relatively simple data distribution, like TOURISM-Monthly. Both autoregressive and non-autoregressive decoding schemes show comparable performance in short-term forecasting, as discussed in the main paper.
@@
D.5 Experiments on Synthetic Datasets
@@ D.7 Model Efficiency
D.8 Further discussion on uni- / multi-variate modeling
Our experiments indicate that the choice between univariate and multivariate modeling is not a primary factor in fulfilling the essential forecasting needs considered in this paper.
D.8.1 Discussion
We discuss the differences between univariate and multivariate modeling from two perspectives:
Dataset Perspective: Whether the dataset is prepared for univariate or multivariate benchmarking.
Model Perspective: How the model handles multivariate data, treating each variable channel independently or not.
Dataset Perspective
All datasets listed in Table 5 are typically referred to as multivariate datasets, indicating that there may be strong connections across different variables. Despite this implication, when developing forecasting models, we can treat each variable channel independently, essentially turning a multivariate dataset into a univariate setup. In contrast, some datasets, like M4, M5, and TOURISM listed in Table 20, explicitly serve univariate modeling. We rarely see multivariate models being developed for these univariate cases.
Model Perspective
We have observed different preferences for univariate and multivariate modeling. Existing models can be categorized into three groups:
Native Univariate Models
Classical models like N-BEATS and N-HiTS.
Most time-series foundation models, such as TimesFM and Chronos.
Native Multivariate Models
Most probabilistic models, such as CSDI and TimeGrad.
Some point forecasting models, such as Informer and Autoformer.
Hybrid Models of Univariate and Multivariate Modeling
Some classical models, such as PatchTST.
Some time-series foundation models, such as MOIRAI.
Native univariate models can also be applied to multivariate datasets by treating them as univariate cases. Similarly, native multivariate models can be applied to univariate datasets by setting the variable dimension to 1. Hybrid models typically include specific modes to activate univariate and multivariate functionalities. For example, in PatchTST, we can use a shared forecasting head for univariate modeling or assign a specific forecasting head for each variable channel to differentiate different variables.
@@ D.8.1 Discussion
D.8.2 Additional Experiments on uni- / multi-variate modeling
In Table 27, we include additional experiments comparing univariate and multivariate modeling of PatchTST across different datasets. Our observation is that there is no definitive answer as to which approach is superior; it depends on the nature of the dataset. Some datasets benefit from modeling variable correlations, while others perform better with independent modeling. The performance gaps are not significant.
Similar observations have been reported in MOIRAI, which allows either univariate or multivariate modes by controlling its cross-variate attention masks. When applied to a downstream forecasting scenario, it can search over the validation set to determine which configurations to activate. We believe such a design could serve as a good example of unifying univariate and multvariate.
Dataset | Pred. Horizon | PatchTST (Multivariate) | PatchTST (Univariate) |
ETTh1 | 96 | 0.3239 | 0.3212 |
192 | 0.3609 | 0.3562 | |
336 | 0.3763 | 0.3737 | |
720 | 0.3882 | 0.3909 | |
ETTm1 | 96 | 0.2652 | 0.2739 |
192 | 0.2926 | 0.2961 | |
336 | 0.3101 | 0.3188 | |
720 | 0.345 | 0.3463 | |
Electricity | 96 | 0.0832 | 0.0857 |
192 | 0.0899 | 0.0912 | |
336 | 0.0995 | 0.1001 | |
720 | 0.1183 | 0.116 | |
Exchange | 96 | 0.0243 | 0.0235 |
192 | 0.0348 | 0.0336 | |
336 | 0.0471 | 0.0462 | |
720 | 0.0787 | 0.0777 | |
Weather | 96 | 0.0872 | 0.0837 |
192 | 0.0924 | 0.0858 | |
336 | 0.0934 | 0.0903 | |
720 | 0.0993 | 0.0953 |
Checklist
1.
For all authors…
Did you describe the limitations of your work? [Yes] See Sections 5.
2.
If you are including theoretical results…
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
4.
If your work uses existing assets, did you cite the creators? [Yes]
Did you mention the license of the assets? [Yes] See Appendix B.4.
5.
If you used crowdsourcing or conducted research with human subjects…